# ChronoFlow: A Data-Driven Model for Gyrochronology

### v1

@@ 3 Age inference framework
We choose to separate the likelihood from the priors in this way so that our model learns the $P_{rot}$ distribution conditioned on the other parameters, and not the overall $P_{rot}$ distribution. By doing so, we mitigate the influence of selection effects in color and age, and so minimize the impact of observational biases on our age inference.
@@
To build the $\mathcal{P}(C_{0}\;|\;\sigma_{C_{0}},\tau)$ prior, we firstly assume that the observed colour $C_{0}$ does not depend on the age of a star. While this is not strictly true (the bluest stars turn off the MS relatively quickly and therefore are more likely to be younger), our model is designed to optimize the conditional distribution of rotation period, so we ignore this effect and assume:
@@
We incorporate $\sigma_{C_{0}}$ by convolving (7) with a Gaussian of variance $\sigma_{C_{0}}^{2}$, so that the functional form of our prior is:
@@
which effectively sums the two probability distributions on the right. Figure 6 illustrates Eqn. (8) as the change in our $C_{0}$ prior as a function of $\sigma_{C_{0}}$. As expected, when $\sigma_{C_{0}}$ is small the uniform $C_{0}$ distribution dominates, but as $\sigma_{C_{0}}$ increases the prior becomes more Gaussian.
@@ 4 ChronoFlow
We trained ChronoFlow to learn $\mathcal{P}(P_{rot}\;|\;C_{0},\sigma_{C_{0}},\tau)$ using the zuko [^4] https://zuko.readthedocs.io/stable/ framework, a python-based package that implements CNFs using the PyTorch [^5] https://pytorch.org/ deep learning library. In particular, we used a Neural Spline Flow (Durkan et al., 2019) with three transform layers and eight bins, trained for 5,000 epochs. We varied our learning rate using a combination of exponential decay (calibrated to decrease from 1e-3 to 5e-7 over 5,000 steps) and a cosine annealing function with a 1,000-epoch period (see Loshchilov & Hutter 2017). This setup was the result of optimizing the model architecture and hyperparameters, such as learning rate and number of transform layers.
@@
Here, $p_{cl}$ is the cluster membership probability for each star (as described in §2), and $p_{out}$ is a universal “outlier probability” of 0.05. $p_{out}$ represents the probability of cluster member straying from the single star rotational sequence (e.g., a blue straggler or an unresolved binary). In future phases, applying a dynamic outlier probability per star that uses indications of binarity (such as radial velocity) could further improve this model.
@@ 5 Results
Figures 7 and 8 present the conditional probability densities calculated by ChronoFlow at the fiducial ages of each of our clusters, including the location of each cluster members in $P_{rot}-C_{0}$ space for reference. An important note is that the “high” probabilities in regions of sparse data at the color extremes are due to the conditional nature of our model. Since each subplot represents $p_{f}(P_{rot}\;|\;C\!,\sigma_{C_{0}},\!\tau)$ at a single age, the probability distribution along any vertical line will integrate to 1. Hence, these “wings” represent the $p_{f}(P_{rot})$ distribution that we would expect to see given stars at those ages and colors, not where we would expect to see a high density of stars overall. This illustrates the importance of applying a conditional normalizing flow to mitigate selection effects, which otherwise could bias our NF in color space.
@@
Figure 11: Inferred cluster ages from LOOCV compared to the fiducial values from literature. Subgroups of the Taurus and Sco-Cen regions are plotted in grey, and the five anomalous clusters are plotted in red. The best fit line is $\tau_{\mathrm{inf}}=(1.01\pm 0.02)\cdot\tau_{\mathrm{fid}}-(0.03\pm 0.06)$, with an intrinsic scatter of 0.06 in $\tau_{\mathrm{inf}}$. This indicates that there is no systematic offset in $\tau_{\mathrm{inf}}$ resulting from the model itself. Error bars include both the statistical uncertainty for each cluster estimate and the intrinsic scatter.
Figure 12: A comparison of the five anomalous clusters from Figure 11. Observed data are plotted over our model’s probability density at both fiducial (top) and inferred (bottom) cluster ages. The same default parameters for $\sigma_{C_{0}}$ and $p_{cl}$ were used as in Figure 7. We exclude stars having $(BP-RP)_{0}$ 0.3 from our age inference, and have shaded this region in red.
@@ 5.1 Stellar and cluster age recovery
Evaluating Eq. (5) over a grid of ages, for a constant set of observed parameters, is straightforward, and this generates a posterior age probability for each star. Figure 10 presents an example of this. As expected, given the almost uniform prior, the posterior follows the likelihood closely.
To obtain the posterior age estimation for a cluster $\mathcal{P}(\tau_{cl})$, we take the product of the individual stellar posteriors for all $n_{cl}$ stars in the cluster:
$\mathcal{P}(\tau_{cl})=\prod_{i=1}^{n_{cl}}\mathcal{P}(\tau_{i}\;|\;P_{rot,i},C_{0,i},\sigma_{C_{0},i})$ (12)
When calculating these cluster posteriors, we exclude stars having $C_{0}$ 0.3 because the stability of our model suffers in this regime. The reason for this behavior is likely twofold: (i) we have sparse sampling; and (ii) those stars may predominantly be above the Kraft break (Kraft 1967), meaning they have radiative envelopes and do not undergo magnetic braking. Beyer & White (2024) identify the Kraft break at $(BP-RP)_{0}\approx 0.54-0.60$, indicating that $C_{0}=(BP-RP)_{0}=0.3$ is a conservative cut and should not exclude any useful data.
@@ 5.1.1 Outliers
We investigated the five anomalous clusters further. Figure 12 presents the observed data for those clusters plotted against our flow results at: (i) the fiducial literature age, and (ii) the age that the final version of ChronoFlow (from which we exclude these five clusters during training) infers. It is clear that for H Persei, the model prediction at the fiducial age is a better visual fit than our inferred age. So, we consider this an outlier that our model is not able to characterize well. In contrast, the other four clusters generally appear to be better visual fits to our inferred ages. We therefore present our inferred ages for these four clusters as new cluster age estimates:
  • M34: $241^{+39}_{-33}$ Myr
  • NGC 2516: $119^{+21}_{-18}$ Myr
  • NGC 1647: $84^{+24}_{-16}$ Myr
  • NGC 1750: $100^{+31}_{-21}$ Myr
These uncertainties include both the statistical uncertainty for each cluster and the intrinsic scatter of our model.
@@ 5.2 Age recovery for single stars
Our focus to this point has been on characterizing the rotational evolution of a population of stars, but now we examine individual stars. As a standalone test, we applied ChronoFlow to the Sun using $P_{rot}$ = 26.09 days and $(BP-RP)_{0}$ = 0.817 (from Popinchalk et al. 2021), and calculated its age to be $5.1^{+1.7}_{-1.4}$ Gyr. This is consistent at $1\sigma$ with the true age of 4.6 Gyr.
@@ 6.1 Other considerations
Another attribute that may affect rotational evolution is metallicity. It would be straightforward to include metallicity as a training parameter in our model, however most of the open clusters are approximately solar metallicity since they are nearby. Given this lack of parameter space coverage combined with the uncertainties on measurements, we do not expect our model to learn anything meaningful from including metallicity. Should significant rotation data become available for stars with super- or sub- solar metallicity, this would be worth investigating.
Another consideration is how to optimize this framework for the lowest mass (e.g. K and M) stars. In that regime, we would expect the Gaia $(G-RP)_{0}$ color to be more descriptive than $(BP-RP)_{0}$ given the low flux in the BP band, so including both as training parameters could improve performance.

### v2

@@ 3 Age inference framework
We choose to separate the likelihood into two terms in this way so that our model learns the $P_{rot}$ distribution conditioned on the other parameters, and not the overall $P_{rot}$ distribution. By doing so, we mitigate the influence of selection effects in color and age, and so minimize the impact of observational biases on our age inference.
@@
To build the $\mathcal{P}(C_{0}\;|\;\sigma_{C_{0}},\tau)$ term, we firstly assume that the observed colour $C_{0}$ does not depend on the age of a star. While this is not strictly true (the bluest stars turn off the MS relatively quickly and therefore are more likely to be younger), our model is designed to optimize the conditional distribution of rotation period, so we ignore this effect and assume:
@@
We incorporate $\sigma_{C_{0}}$ by convolving (7) with a Gaussian of variance $\sigma_{C_{0}}^{2}$, so that the functional form becomes:
@@
which effectively sums the two probability distributions on the right. Figure 6 illustrates Eqn. (8) as the change in $C_{0}$ as a function of $\sigma_{C_{0}}$. As expected, when $\sigma_{C_{0}}$ is small the uniform $C_{0}$ distribution dominates, but as $\sigma_{C_{0}}$ increases then $\mathcal{P}(C_{0}\;|\;\tau,\sigma_{C_{0}})$ becomes more Gaussian.
@@ 4 ChronoFlow
We trained ChronoFlow to learn $\mathcal{P}(P_{rot}\;|\;C_{0},\sigma_{C_{0}},\tau)$ using the zuko [^4] https://zuko.readthedocs.io/stable/ framework, a python-based package that implements CNFs using the PyTorch [^5] https://pytorch.org/ deep learning library. In particular, we used a Neural Spline Flow (Durkan et al., 2019) with three transform layers and eight bins, trained for 5,000 epochs. We varied our learning rate using a combination of exponential decay (calibrated to decrease from 1e-3 to 5e-7 over 5,000 steps) and a cosine annealing function with a 1,000-epoch period (see Loshchilov & Hutter 2017). We chose this setup after experimenting with a variety of model architectures, learning rates, and number of transform layers, as it resulted in the least test loss (i.e., was able to recover our cluster ages with the highest accuracy, using leave-one-out cross validations tests as described in §5.1).
@@
Here, $p_{cl}$ is the cluster membership probability for each star (as described in §2), and $p_{out}$ is a universal “outlier probability” of 0.05. $p_{out}$ represents the probability of cluster member straying from the single star rotational sequence (e.g., a blue straggler or an unresolved binary). In future phases, applying a dynamic outlier probability per star that uses indications of binarity (such as radial velocity) could further improve this model (see §6.1 for further discussion of binarity).
@@ 5 Results
Figures 7 and 8 present the conditional probability densities calculated by ChronoFlow at the fiducial ages of each of our clusters, including the location of each cluster members in $P_{rot}-C_{0}$ space for reference. An important note is that the “high” probabilities in regions of sparse data at the color extremes are due to the conditional nature of our model. Since each subplot represents $p_{f}(P_{rot}\;|\;C_{0}\!,\sigma_{C_{0}},\!\tau)$ at a single age, the probability distribution along any vertical line will integrate to 1. Hence, these “wings” represent the $p_{f}(P_{rot})$ distribution that we would expect to see given stars at those ages and colors, not where we would expect to see a high density of stars overall. This illustrates the importance of applying a conditional normalizing flow to mitigate selection effects, which otherwise could bias our NF in color space.
@@
Figure 11: Inferred cluster ages from LOOCV compared to the fiducial values from literature. Subgroups of the Taurus and Sco-Cen regions are plotted in grey, and the five anomalous clusters are plotted in red. The best fit line is $\tau_{\mathrm{inf}}=(1.00\pm 0.02)\cdot\tau_{\mathrm{fid}}-(0.00\pm 0.06)$, with an intrinsic scatter of 0.06 in $\tau_{\mathrm{inf}}$. This indicates that there is no systematic offset in $\tau_{\mathrm{inf}}$ resulting from the model itself. Error bars include both the statistical uncertainty for each cluster estimate and the intrinsic scatter.
Figure 12: A comparison of the five anomalous clusters from Figure 11. Observed data are plotted over our model’s probability density at both fiducial (top) and inferred (bottom) cluster ages. The same default parameters for $\sigma_{C_{0}}$ and $p_{cl}$ were used as in Figure 7. We exclude stars having $(BP-RP)_{0}<$ 0.5 from our age inference, and have shaded this region in red. The ChronoFlow estimates for M34 and NGC 2516 are consistent with literature age estimates of M34 (Gaia Collaboration et al., 2018; Ianna & Schlemmer, 1993; Meibom et al., 2011) and NGC 2516 (Meynet et al., 1993; Godoy-Rivera et al., 2021; Sung et al., 2002; Richer et al., 2021; Bouma et al., 2021).
@@ 5.1 Stellar and cluster age recovery
Evaluating Eq. (5) over a grid of ages, for a constant set of observed parameters, is straightforward, and this generates a posterior age probability for each star. Figure 10 presents an example of this. As expected, given the uniform prior, the posterior follows the likelihood closely.
To obtain the posterior age estimation for a cluster $\mathcal{P}(\tau_{cl})$, we take the product of the individual stellar likelihood functions for all $n_{cl}$ stars in the cluster, and multiply that by our age prior $\mathcal{P}(\tau\;|\;\sigma_{C_{0}})$. Since the stellar likelihood consists of both the $\mathcal{P}(P_{rot}\;|\;C_{0},\sigma_{C_{0}},\tau)$ and $\mathcal{P}(C_{0}\;|\;\tau,\sigma_{C_{0}})$ terms, we calculate each cluster posterior using:
$\mathcal{P}(\tau_{cl})=\mathcal{P}($ $\tau\;|\;\sigma_{C_{0}})$ (12)
$\cdot\prod_{i=1}^{n_{cl}}[\mathcal{P}$ $(P_{rot,i}\;|\;C_{0,i},\sigma_{C_{0},i},\tau)$
$\cdot\mathcal{P}(C_{0,i}\;|\;\tau,\sigma_{C_{0},i})]$
We have ignored the evidence term here because we re-normalize the cluster posteriors to integrate to 1.666Functionally, since we have applied a uniform age prior and are normalizing $\mathcal{P}(\tau_{cl})$, we can instead calculate the cluster posterior as the product of all stellar posteriors in the cluster. Figure 10 illustrates the cluster posterior calculation for NGC 2281 as an example.
When calculating these cluster posteriors, we exclude stars having $C_{0}>$ 0.5 because the stability of our model suffers in this regime. The reason for this behavior is likely twofold: (i) we have sparse sampling; and (ii) those stars may predominantly be above the Kraft break (Kraft 1967), meaning they have radiative envelopes and do not undergo magnetic braking. Beyer & White (2024) identify the Kraft break at $(BP-RP)_{0}\approx 0.54-0.60$, indicating that $C_{0}=(BP-RP)_{0}=0.5$ is a reasonably conservative cut that should not exclude any useful data.
@@ 5.1.1 Outliers
Here we discuss possible reasons for the poor match between observations and ChronoFlow for the aforementioned five anomalous clusters.
H Persei has an abundance of fast rotators relative to other clusters of similar ages, and a more dispersed slowly rotating sequence. Moraux et al. (2013) note that the fast rotators may represent a population of tidally locked close binary systems. Getman et al. (2023) suggest that the $P_{rot}$ measurements may also be contaminated by aliases from solar and lunar cycles. This possible overabundance of fast rotators does not intuitively explain why our model over-estimates the age of H Persei, but these factors suggest that the observations we are using may not be an accurate representation of the $P_{rot}$ evolution of single stars at that age.
Our inferred age for M34 of $245^{+40}_{-34}$ Myr does not match the ages from Cantat-Gaudin et al. 2020 (132 Myr) or Bossini et al. 2019 (126 Myr) well, but is more consistent with Gaia Collaboration et al. 2018 (251 Myr). There is also significant variation in other literature isochronal ages, ranging from e.g. $132\pm 63$ Myr (Richer et al. 2021) to 250 Myr (Ianna & Schlemmer 1993). Notably, James et al. (2010) and Meibom et al. (2011) inferred rotational ages of $193\pm 9$ Myr and 240 Myr respectively, so our age is consistent with this larger set of literature estimates.
NGC 2516 is another cluster for which literature values vary significantly. The average isochronal age from our main catalogs is 263 Myr, but the next oldest literature age is 200 Myr (Richer et al. 2021), and we find four more even younger ages (the youngest being $102\pm 15$ Myr from Li et al. 2024). Our estimate of $132^{+20}_{-18}$ Myr is consistent with these younger literature values , and also consistent with the rotational ages estimated by Godoy-Rivera et al. 2021 (150 Myr).
We only have 14 stars in NGC 1750 having $C_{0}>$ 0.5, and the reddest has $C_{0}$ = 1.25, so the small sample size and lack of color coverage may hinder our age inference for this cluster. NGC 1750 is also less studied than some of the other clusters, with only two isochronal ages we could find outside of our main catalogs. These two estimates are younger than the fiducial age of 323 Myr ($257\pm 123$ Myr from Richer et al. 2021 and $200\pm 50$ Myr from Galadi-Enriquez et al. 1998), but still not consistent with our inferred age of $99^{+32}_{-21}$ Myr.
Small sample size and color coverage may also affect NGC 1647, for which we have 19 total stars with $C_{0}>0.5$ and a maximum $C_{0}$ of 1.26. More isochronal age estimates are available in literature for this cluster, and similar to NGC 1750 they are on average younger than those from our main catalogs (fiducial age of 309 Myr), with five ages ranging from 110 (Turner, 1992) to 200 (±$\pm 50$; Richer et al. 2021) Myr. Again, these are closer to (but still inconsistent with) our inferred age of $87^{+25}_{-18}$ Myr.
Figure 12 presents the observed data for those clusters plotted against our flow results at: (i) the fiducial literature age, and (ii) the age that the final version of ChronoFlow (from which we exclude these five clusters during training) infers. It is clear that H Persei seems to be an outlier than our model is unable to characterize well. Our model fits NGC 1647 and NGC 1750 better at our inferred ages than the fiducial ages, so we present that as tentative evidence that these clusters are on the younger end of literature estimates, however we do not claim these age estimates to be accurate due to the lack of sample size and color coverage. In contrast, M34 and NGC 2516 are well sampled, and our inferred ages fit the observations, so we present these as well-motivated rotational ages:
  • M34: $245^{+40}_{-34}$ Myr
  • NGC 2516: $132^{+20}_{-18}$ Myr
which support literature estimates of M34 (Gaia Collaboration et al., 2018; Ianna & Schlemmer, 1993; Meibom et al., 2011) and NGC 2516 (Meynet et al., 1993; Godoy-Rivera et al., 2021; Sung et al., 2002; Richer et al., 2021; Bouma et al., 2021). Our reported uncertainties include both the statistical uncertainty for each cluster and the intrinsic scatter of our model.
@@ 5.2 Age recovery for single stars
Our focus to this point has been on characterizing the rotational evolution of a population of stars, but now we examine individual stars. As a standalone test, we applied ChronoFlow to the Sun using $P_{rot}$ = 26.09 days and $(BP-RP)_{0}$ = 0.82 (from Casagrande & VandenBerg 2018), and calculated its age to be $5.1^{+1.7}_{-1.4}$ Gyr. This is consistent at $1\sigma$ with the true age of 4.6 Gyr.
Figure 15: Average residuals as a function of color and inferred age. Data are averaged over bins of width 0.2 in both dimensions. Grey shaded regions are bins with no data.
Figure 16: Average standard deviation of individual stellar posteriors as a function of color and inferred age. Data are averaged over bins of width 0.2 in both dimensions. Grey shaded regions are bins with no data.
@@
Therefore, while we summarize the statistical uncertainty for individual stars as $\approx$ 0.7 dex, these systematics should be considered when using ChronoFlow for individual stellar age estimates.
It is also important to note that there is almost certainly some level of contamination in our cluster sample, and therefore for non-members we are incorrectly using the cluster ages as the true ages of these stars. However, this effect is difficult to quantify since by definition the contaminants have not been identified.
5.3 Comparison to literature models
Here we discuss the results and performance of ChronoFlow relative to the most comparable literature models: GPgyro [^8] https://github.com/lyx12311/GPgyro by Lu et al. (2024a) and gyro-interp [^9] https://github.com/lgbouma/gyro-interp by Bouma et al. (2023). We performed two tests: (i) a cluster age recovery exercise, and (ii) a residual analysis analogous to §5.2.
Since the other two models use $T_{\text{eff}}$ instead of $(BP-RP)_{0}$, we restricted the stars used in these tests to the subset of our catalog having a measured Gaia DR3 teff_gspphot. Additionally, we follow the guidance of the other models’ effective parameter space to perform additional cuts. For GPgyro, we apply:
  • 3000 $\leq$ teff_gspphot $\leq$ 7000
  • Absolute Gaia DR3 magnitude $M_{G}<4.2$ (derived from DR3 parallax)
  • Stellar age (from cluster membership) $\geq$ 670 Myr
It is also worth noting that GPgyro is not expected to perform as well for stars $\lesssim$ 1.5 Gyr, however we still include these for completeness. For gyro-interp, we apply:
  • 3800 $\leq$ teff_gspphot $\leq$ 6200
  • Stellar age (from cluster membership) $\geq$ 80 Myr
For the cluster age recovery exercise, we first computed the stellar posteriors with each model. Since GPgyro outputs stellar median ages with uncertainties, and not an age probability distribution, we computed the cluster estimates as the distribution of the median inferred stellar ages. While gyro-interp outputs complete stellar age posteriors, since it does not include an outlier model that forces a non-zero probability at every age, taking the product of stellar posteriors directly often results in a null cluster posterior. So, we calculated cluster age probabilities in two ways. The first method was the same as for GPgyro. In the second, we manually added the same 5% outlier probability that we include in ChronoFlow to the individual gyro-interp stellar posteriors, and then multiplied them to get the cluster posteriors. Both results are presented in Figure 17.
It should also be noted that the datasets that we have used to compare the models likely introduce several systematics in the cluster age recovery test. First, the other models were not trained on Gaia DR3 $T_{\text{eff}}$, so age inference using that temperature prescription may decrease the performance. However, we considered modeling $T_{\text{eff}}$ for our datalog to be outside the scope of this study. A competing effect is that we used LOOCV when testing the cluster age recovery of ChronoFlow in this context, but our data catalog includes stars that were used to train the other models, so we expect their performance to be better given that the test data is not entirely new. We have not attempted to quantify these effects, but to first order assume that they cancel out and ignore them in our evaluation.
The results of the age inference test (Figure: 17) demonstrate that in the region of parameter space where the models overlap, the performance across models is comparable. The benefit of ChronoFlow is that it extends to younger ages than the other two. We also see that training ChronoFlow on this subset of data decreases performance significantly compared to the model trained on our entire data catalog, indicating that the extent of our training catalog is an an important factor in performance, and perhaps that additional systematics are introduced by only considering stars with $T_{\text{eff}}$ measurements in Gaia DR3.
Figure 18 presents the residual analysis for all three models (the left panel is a copy of Figure 15). It is apparent that the applicable parameter space of ChronoFlow extends younger than that of the other two models, both overall and when considering regions of low bias. There is also wider coverage across the color space, although the other models use $T_{\text{eff}}$ in place of color so that comparison is more difficult. In the regime of overlap between ChronoFLow and GPgyro, the models are comparable and both relatively unbiased. GPgyro is more biased when applied at lower inferred ages, however applying the stricter cut of $>1.5Gyr$ to the test data set removes most of that bias (although also reduces the parameter space; see the farthest right panel). Where ChronoFlow and gyro-interp overlap, both models perform well, however the residuals are systematically slightly higher in gyro-interp. This indicates that ChronoFlow as a model may impose less structure across the parameter space, and therefore demonstrate better flexibility in age inference.
Figure 17: Cluster age inference results from each model when applied to the subset of our data catalog in the models’ effective regimes. The outlier clusters described in §5.1.1 are included. Only stars with a Gaia DR3 teff_gspphot measurement were included in this exercise, hence the discrepancy with Figure 11. ChronoFlow was evaluated using LOOCV, while we used the published version of the other models. We shade the 95% credible interval of a linear fit to $\tau_{\text{inf}}$ vs. $\tau_{\text{fid}}$ for each model. Overall, the accuracy of each model is comparable, however ChronoFlow maintains its accuracy to younger ages than the other models.
Figure 18: A comparison of the stellar residuals between ChronoFlow, gyro-interp, and GPgyro in $\tau_{\text{inf}}-(BP-RP)_{0}$ space, averaged across bins of 0.2 in both dimensions. The residual is the inferred age from each model minus the fiducial age according to our catalog. The left panel is a copy of Figure 15. It is evident that ChronoFlow has more complete coverage over the parameter space at open cluster ages than the other models, and it exhibits less to equivalent systematic uncertainty compared to the other models in the regions of overlap.
@@ 6 Systematic Uncertainties
6.1 Impact of binarity
We have treated potential binary contamination simplistically in this work, by first applying cuts in an attempt to restrict our training data to only single stars (using Gaia DR3 RUWE and literature flags), and then including a uniform outlier component in our model to account for binaries that passed those cuts. We recognize that this uniform outlier is naive; there are additional data such as radial velocities that we could use to better inform the outlier probability for each star, and it would also be possible to model a binary evolution track in ChronoFlow in addition to the single star track. This could help account for systematic rotational trends, such as the overdensity of tidally synchronized binaries among fast rotators in the Kepler field identified by Simonian et al. (2019), and the abundance of binaries in the fast rotating sequence of Blanco 1 and the Pleiades found by Gillen et al. (2020). Intentionally modeling binaries would also be preferable to attempting to remove them entirely from our model, since (i) there will likely always be some level of contamination, and (ii) binaries make up a significant portion of the stellar population. The behavior of binary contaminants may therefore be an additional unmitigated source of systematics in our current model, but we leave a more comprehensive treatment for future work.
@@ 6.2 Other considerations
While we have mitigated the effects of detection bias in color and age by using a conditional normalizing flow, selection effects in $P_{rot}$ still exist. $P_{rot}$ measurements are naturally biased towards stars with shorter periods (since these rotate more times per observational baseline), and greater periodic variability amplitude (which results in a stronger signal). Overall, the concentration of cluster members at reliably measurable periods indicates that we do not expect there to be a significant population of stars at undetectable periods, at least up to the age of M67 where we do see some very long rotators. However, older fainter stars (or high mass stars) with weak magnetic activity may have low rotational modulation amplitudes due to a lack of star spots, and young stars may have extreme short term activity-induced variability that obscures the rotational signal. In both cases, our ability to measure a rotation period is hindered. We leave that as a caveat to this work, but it also begs the question of whether including activity-based metrics in a model such as ChronoFlow can improve age constraints; this is a topic we intend to investigate in future work.
Another attribute that affects rotational evolution is metallicity, as its influence on opacity affects the depth of the surface convective zone and therefore the amount of magnetic braking (e.g. Claytor et al. 2020; Lu et al. 2024a). It would be straightforward to include metallicity as a training parameter in our model, however most of the open clusters are approximately solar metallicity since they are nearby. Given this lack of parameter space coverage combined with the uncertainties on measurements, we do not expect our model to learn anything meaningful from including metallicity. Should significant rotation data become available for stars with super- or sub- solar metallicity, this would be worth investigating.
In future work, it is important to consider how to optimize this framework for the lowest mass (e.g. K and M) stars. In that regime, the Gaia $(BP-RP)_{0}$ is not as descriptive given the low flux in the BP band. This is seen for M67 in Figure 7; there appear to be many stars above the slow rotator sequence in a narrow color range, but when Dungee et al. (2022) plot those rotation periods as a function of stellar effective temperature ($T_{\text{eff}}$) the sequence is smooth. We would expect the Gaia $(G-RP)_{0}$ color to be more descriptive than $(BP-RP)_{0}$, so including both as model parameters could improve performance.